OpenStudy (anonymous):

What is 1 2/3 + 3 4/6

OpenStudy (anonymous):

i think its 6.3 :)

OpenStudy (anonymous):

5/3 + 22/6 = 10/6 + 22/6 = 32/6 = 5 2/6 = 5 1/3

OpenStudy (anonymous):

5/3 + 22/6 =1.6666+3.6666=5.333